# secReferences in the Online Appendix

### v4

@@ 1 Introduction
This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value. We focus on campaign timing—one of the few observable aspects of NGO activity, which remains far more opaque than that of for-profit firms. [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base. Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored. While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior. We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
@@
Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize visibility by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals. However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly. Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
@@
Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008). However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns. In contrast, past non-AGM campaigns have no influence on proposal outcomes. NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
@@
Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time. This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant. Our main endogenous variable, therefore, is NGO campaign timing, which we explain through NGO experience: more experienced NGOs campaign earlier, trading visibility for impact. While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience. This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
@@ 2 Data
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
@@
We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. [^3] Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
@@ 3 AGM Dates
Target selection. Our analysis focuses on campaign timing, abstracting from NGOs’ selection of a target, which we leave for future research. However, Appendix Table A1 shows that NGOs’ targets does not depend on prior campaign experience, a key variable of our analysis. Panel (c) finds no differences in market capitalization, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs. Panel (d) regresses firm characteristics on NGOs’ past campaigns, showing no correlation between target selection and NGO experience. However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
@@ 6.1 Accumulating Visibility
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
@@ 7 Discussion and Conclusion
By analyzing the timing of NGO campaigns, we demonstrate that this seemingly minor decision has substantial consequences for both targeted firms and the effectiveness of future initiatives. Furthermore, we demonstrate that NGOs’ tactics evolve over time as they seek to either indirectly pressure firms (“exposure”) or directly shape their operations (“influence”). Thus, systematically analyzing NGOs’ behavior can shed new light on how activism helps firms internalize externalities.
@@
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can discipline managers (Levit and Malenko, 2011). Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results. Also, we provide evidence suggesting that NGO campaigns on AGMs do not target different firms compared to campaigns at other times, which helps mitigate selection issues (Appendix Table A2).
@@
Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics. Indeed, we find that mature NGOs—those with multiple prior campaigns in the sample—no longer inflict additional reputation losses by campaigning on the AGM date, whereas they did so in their first campaigns (compare the blue dots and red triangles in Appendix Figure B4). In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex, with multiple ways to influence firms) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
@@ Appendix B Omitted Figures
Figure B8: Examples of positive and negative E&S news

### v5

@@ 1 Introduction
This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value. We focus on campaign timing—one of the few observable aspects of NGO activity, which is far more opaque than that of for-profit firms. [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base. Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored, especially outside developing-countries contexts (e.g., Deserranno et al., 2025). While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior. We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
@@
Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize exposure by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals. However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly. Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
@@
Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008). However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns. In contrast, past non-AGM campaigns have no influence on proposal success. NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
@@
Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time. This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant. Our main endogenous variable, therefore, is NGO campaign timing, which we explain by reference to each NGO’s prior campaign experience: more experienced NGOs campaign earlier, trading visibility for impact. While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience or campaign timing. This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
@@ 2 Data
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. Those single-firm campaigns represent 93% of the original sample. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
@@
We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. [^3] Evidence shows that shareholders generally agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
@@ 3 AGM Dates
Target selection. While our analysis centers on campaign timing—leaving NGOs’ target selection to future research—Appendix Table A2 shows that NGOs campaigning on AGM dates do not select firms that differ in observable characteristics (market capitalization, assets, sales, media coverage, or environmental and social scores) from those targeted at other times, and campaign sentiment is likewise comparable across both groups. Moreover, Table A1 demonstrates that target choice is independent of prior campaign experience: Panel (c) reveals no significant differences in market cap, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs, and Panel (d) finds no correlation between firm characteristics and NGOs’ past campaigns. However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
@@ 6.1 Accumulating Visibility
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of shareholder proposals.
@@ 7 Discussion and Conclusion
Examining NGO campaign timing reveals that not just the decision to intervene, but its timing, can materially affect firm behavior and shape the course of future activism. Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
@@
However, campaigning before the AGM offers direct advantages. Specifically, when an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of acceptance—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3). Given that proposals typically attract low support, early NGO interventions can effectively discipline managers (Levit and Malenko, 2011). Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results. Furthermore, we show that selection issues are not a central concern: Panel d of Appendix Table A1 shows target‐firm characteristics are invariant to an NGO’s past campaign experience, and Appendix Table A2 confirms NGOs campaigning at AGMs do not systematically target different firms than those campaigning at other times.
@@
Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics. Consistent with this, mature NGOs—those with multiple prior campaigns—do not inflict greater reputation harm by campaigning on the AGM date, whereas they do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B4). In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
@@ Appendix B Omitted Figures
Figure B8: NGO campaigns and consumer backlash